OpenStudy (anonymous):

6k-3k2)=3(k-2) step by step plz

OpenStudy (anonymous):

I dont get it because theres no opening to the first parenthisis> )

OpenStudy (anonymous):

is this suppose to read as 6k-3k^2=3(k-2)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay. 6k-3(k+2)=3(k-2) multiply -3 by each term inside the parentheses 6k-3k-6=3(k-2) since 6k and -3k are ike terms, add -3k to 6k to get 3k 3k-6=3(k-2) multiply 3 by each term inside the parentheses again 3k-6=3k-6 since 3k has the variable to solve for, move it to the left hand side of the equation by subtracting 3k from both sides 3k and -3k are like terms add -3k to 3k to get o 0-6=-6 remove the 0 from the polynomial; adding or subtracting 0 doesn't change the value of the expression. -6=-6 since -6=-6, the expression will always be true.